In 2017, a relevant benchmark for FIN corporation is found to have the following values.

P/E = 8.1

P/S = 0.75

P/B = 2.5

P/OCF = 6.0

Calculate the estimated price for FIN corp using the valuation by comparables.

(FIN corp Balance Sheet)

2016 2017
Cash 499 626
Receivables 3831 4141
Inventories 4449 4809
Other current assets 869 869
Total current assets 9648 10445
Property plant and equipment PPE 18442 21861
accumulated depreciation 4909 6172
Net property, plant and equipment 13533 15689
Other concurrent assets 973 973
Total assets 24154 27107
Accounts payable 4160 4496
Current portion of long term debt 905 892
Accrued expenses 1566 1693
Income taxes payable 423 424
Total current liabilities 7054 7505
Defferred income taxes and other liabilities 1152 1152
Long term Debt 8088 9396
Total Liabilities 16294 18053
Preferred stock 0 0
Common Stock 75 75
Capital surplus.    1098 1118
Retained Earnings.    6687 7861
Shareholders equity 7860 9054
Total liabilities and net worth 24154 27107

(Fin corp Income statement.)

2017

Net Sales 43115
Cost of goods sold 29450
Gross profit 13665
Selling, general, and other administrative expenses 9602
Depreciation and amortization expense 1263
interest expense 578
Income before taxes 2222
Income taxes expense 844
Net income 1378
Outstanding shares 905
share price $25
Answer #1

Using

1)

P/E = 8.1

EPS for FIN = Net Income/ Total shares o/s = 1378/905 = 1.52

Price =8.1 * 1.52 = $ 12.33

2)

P/S = 0.75

Sales for FIN = 43115

Sales/share = 43115/905 = $ 47.64

Price = 0.75 * 47.64 = $ 35.73

3)

P/B = 2.5 times

BV per share = (Total Assets -Liabilities)/ Number of shares = (27107 - 18053)/905 = 10.004

Price = 2.5 * 10.004 =$ 25.01